OpenStudy (lovelyanna):

Which is the equation of the given line in point-slope form? A. y + 1 = 6/7 (x + 4) B. y + 1 = 6/11 (x + 4) C. y + 4 =11/6 (x + 1) D. y - 1 = 6/11 (x - 4)

OpenStudy (lovelyanna):

@jigglypuff314

jigglypuff314 (jigglypuff314):

well... idk the given line ;) point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1)

OpenStudy (lovelyanna):

Hold on, I'm pasting the line :P

jigglypuff314 (jigglypuff314):

first find slope: \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]when given points (x1, y1) and (x2, y2)

OpenStudy (lovelyanna):

I know what slope equals :P

OpenStudy (lovelyanna):

Is the slope of the line -2?

jigglypuff314 (jigglypuff314):

2 - (-1) = 2 + 1 = 3 1.5 - (-4) = 1.5 + 4 = 5.5 so 3/5.5 or 6/11

OpenStudy (lovelyanna):

Oh.. I must have done it wrong.. sorry :(

jigglypuff314 (jigglypuff314):

it's fine :P everyone makes mistakes :)

OpenStudy (lovelyanna):

Thats true :P So now its either B or D, right?

jigglypuff314 (jigglypuff314):

yep :) try plugging in point (-4, -1) into y - y1 = (6/11)(x - x1)

jigglypuff314 (jigglypuff314):

y - y1 = (6/11)(x - x1) y -(-1) = (6/11)(x-(-4)) y + 1 = (6/11)(x + 4)

OpenStudy (lovelyanna):

Ok, now I feel stupid and dumb.. >.>

jigglypuff314 (jigglypuff314):

lol ;) good job! :D

OpenStudy (lovelyanna):

Ok, thanks :) I have 4 left, you think you could help with the rest? :P

jigglypuff314 (jigglypuff314):

yea sure :)

OpenStudy (lovelyanna):

Thanks! :) Which is an equation of the given line in standard form? A. –8x + 9y = 23 B. –8x + 9y = –23 C. –8x + 7y = 25 D. –9x + 8y = –23 http://static.k12.com/calms_media/media/1331000_1331500/1331297/1/c5c8e45fd70b35572ab9141472538d3ecb31705d/HS_AL1G2_S1_07_10_ART2.jpg

jigglypuff314 (jigglypuff314):

I not like standard form >:( first find slope the plug into what ever equation format you like the best then simplify it into standard form...

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Awesome! At least I got one right :P

OpenStudy (lovelyanna):

Which is the equation of the given line? A. x = –3 B. x = 3 C. y = –3 D. y = –3x http://static.k12.com/calms_media/media/1312500_1313000/1312643/1/99a45d74ac85e7c581b85f02ff88f26b08902e7d/HS_A1G2_S1_07_10_ART3.jpg

jigglypuff314 (jigglypuff314):

horizontal line's equation is y = (something) with no x involved

OpenStudy (lovelyanna):

So, find the slope of the line first?

jigglypuff314 (jigglypuff314):

you can see that it's a horizontal line therefore slope = 0 (if it was a vertical line then slope = undefined)

jigglypuff314 (jigglypuff314):

no x involved, because having an x in there would give it slope

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (lovelyanna):

Thanks :) Let me write the next one..

OpenStudy (lovelyanna):

Which is the equation of the given line in slope-intercept form? A. y =3x + 2 B. y +3x - 2 C. y = -3x + 2 D. y = -1/3x + 2 I can't paste the graph but the line goes like this: